OpenStudy (anonymous):

5,000(1.00283)^22 Can I get help simplifying please?:) It's an exponential function and I'm not coming out to the right amount..

OpenStudy (whpalmer4):

You are using a calculator to evaluate this? How exactly are you entering it? What do you think the right amount is?

OpenStudy (anonymous):

I just don't know where to put in the ^22 part.. And I've tried quite a few ways..

OpenStudy (tkhunny):

Do you have a key that looks like this: \(x^{y}\) ?

OpenStudy (anonymous):

No.. I don't think so...

OpenStudy (whpalmer4):

@Mgirl22737 what brand/model of calculator do you have? If you don't have a \(y^x\) key, or anything that looks like it, you've got a couple of options: a) use an online calculator in your web browser, such as http://web2.0calc.com/ b) if you have log functions (\(\log, 10^x, \ln, e^x\)) on your calculator, you could evaluate that by taking the log of 1.00283, multiplying by 22, taking the antilog (use e^x if you used ln, or 10^x if you used log), then multiplying by 5000. c) you could enter 5000*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283*1.00283 d) get a *really* big piece of paper, and do option c) by hand :-)

OpenStudy (anonymous):

It's a texas Intruments TI-85 collage calc.. @whpalmer and thanks:)

OpenStudy (whpalmer4):

Ah, okay, your calculator has a key that looks like ^ on the right side that should do the trick. 1.00283 ^ 22 * 5000 = should do the trick. I last owned a TI calculator in 1982 or so, I think, and it made me draw my own graphs :-)